# Improving SDSS Cosmological Constraints through $\beta$-Skeleton Weighted Correlation Functions

### v1

@@ III.1 $\beta$-skeleton analysis
The deviations between the real data and COLA simulations, regardless of the $\Omega_{m}$ values used, are considerably small, with no significant differences observed. The mean comoving distance $r$ is 1346.1 $h^{-1}{\rm Mpc}$ for CMASS, 1345.7 $h^{-1}{\rm Mpc}$ for $\Omega_{m}=0.25$, 1346.7 $h^{-1}{\rm Mpc}$ for $\Omega_{m}=0.31$, and 1344.8 $h^{-1}{\rm Mpc}$ for $\Omega_{m}=0.4$. These values suggest that when varying $\Omega_{m}$, the changes in $r$ for COLA are on the order of 0.1%, making it challenging to rely solely on this quantity for distinguishing between cosmologies.
Similarly, the density distributions of the connection length $L_{\rm con}$ and the orientation of these connections, $\mu$, are illustrated in the middle and right panels. We observe that the discrepancies in either $L_{\rm con}$ or $\mu$ between real data and COLA simulations are statistically insignificant. For instance, the mean $L_{\rm con}$ is 8.83 ${}h^{-1}{\rm Mpc}$ for CMASS, and 8.70, 8.68, and 8.71 $h^{-1}{\rm Mpc}$ for $\Omega_{m}=0.25,0.31,$ and $0.4$, respectively, leading to deviations of less than $2\%$ across all cases. Additionally, the mean values of $\mu$, averaged over the range of $\mu\in[0,0.97]$, are 0.50 for CMASS, and 0.51, 0.52, 0.53 for $\Omega_{m}=0.25,0.31,0.4$ in COLA, respectively. This suggests that the directions of the connections in all cases are consistently randomly distributed, with no preferred orientation. Moreover, we can clearly observe a rise in density when $\mu>0.9$. This is likely due to the Finger of God (FoG) effect, increasing the orientation along LOS.
Figure 4: Histogram of the number of neighbors around each galaxy, referred to as connectivity ($N_{\rm con}$), in the range of $[0,5]$ within the $\beta$-skeleton cosmic webs. It peaks at $N_{\rm con}=2$ and then decreases rapidly. The variation in connectivity is negligible for both CMASS data and COLA mocks under three different $\Omega_{m}$ values.
@@
On the right side, four violin plots depict the dependence of $\bar{D}_{\rm nei}$ distributions on connectivity $N_{\rm con}$, ranging from 1 to 5. The cases of CMASS-NGC data and COLA mocks under three different $\Omega_{m}$ values are presented sequentially. As seen, all COLA results are nearly identical to the CMASS case when $N_{\rm con}<5$. However, when $N_{\rm con}=5$, the distribution of $\bar{D}_{\rm nei}$ exhibits slight variation, with mean values of $\bar{D}_{\rm nei}=11.52,12.38,11.27,12.78~{}h^{-1}{\rm Mpc}$ for CMASS and COLA simulations with $\Omega_{m}=0.31$, $0.25$ and $0.4$.
@@ III.2 MCFs using the $\beta$-skeleton statistics
Briefly, the procedure for calculating the MCFs remains the same as the standard one; however, the weights are assigned based on the measures of the $\beta$-skeleton statistics. Compared with the traditional 2PCF, defined as $\xi(\bm{r})=\langle\delta(\bm{x})\delta(\bm{x}+\bm{r})\rangle$, the form of MCF, following Yang et al. (2020b), is given by
@@
where the term $w(\bm{x})$ denotes the chosen weights used in MCFs. The term $\delta(\bm{x})$ denotes the point-like density contrast, expressed as $\delta(\bm{x})=\delta{\rho}/\bar{\rho}$.
@@
We utilize the widely adopted Landy–Szalay estimator Landy and Szalay (1993), expressed as:
@@
The random sample comprises 10 times the number of objects compared to the CMASS galaxies, and each PATCHY simulation for covariance estimation includes 20 times the number of objects than CMASS. Based on our testing, results converge with a random sample containing about 10 times more objects than CMASS, and the weights for the random samples are consistently set to unity.
@@ III.3 One-dimensional MCFs: $\hat{W}_{0}(s)$ and $\hat{W}_{\Delta s}(\mu)$
By integrating $W(s,\mu)$ over either $s$ or $\mu$, two one-dimensional statistical quantities can be defined. The first one is the monopole of the MCF, a function of the clustering scale, represented as:
@@
The second quantity is the $\mu$-dependent function, represented as:
@@
To mitigate the impact of galaxy bias and enhance the accuracy of the analysis, it is common to employ normalized statistics, which utilize the shape rather than the amplitude to extract cosmological information. By doing this, the normalized quantities based on Eqs. 4 and 5 can be expressed as follows:
@@
In choosing $a=15~{}h^{-1}{\rm Mpc}$, $b=57~{}h^{-1}{\rm Mpc}$, and $\mu_{\max}=0.97$ (to reduce effects of fiber collisions and RSDs), extensive testing indicates that the selected parameters are effective for studying clustering features and for enhancing the sensitivity on different cosmologies. From tests, we find that opting for equally spaced divisions of $s$ within the range of $s\in[15,57]$ and $\mu$ in the interval $\mu\in[0,0.97]$ into 7 bins results in optimal and robust constraints on $\Omega_{m}$.
@@ IV Results
Figure 6: Comparison of one-dimensional MCFs with different weighting schemes for the CMASS-NGC dataset. Left: In comparison to the 2PCF (red solid), the monopole of MCFs, $s^{2}W_{0}(s)$ (as defined in Eq. 4), is shown for three different weights: $N_{\rm con}$ (blue dashed-dotted), $\bar{D}_{\rm nei}$(orange dashed), and $1/\bar{D}_{\rm nei}$ (green dotted). Right: Same as in the left panel, but for focusing on the anisotropic clustering in MCFs, $W_{\Delta s}(\mu)$ (as defined in Eq. 5). As observed, using the weight $1/\bar{D}_{\rm nei}$ leads to the highest magnitudes in both panels. This is because a number of objects in underdense regions are assigned a larger weight, significantly enhancing the amplitudes of these two-point functions. Conversely, assigning $\bar{D}_{\rm nei}$ as the weight will lead to the lowest amplitudes in both panels. In either the left or right panel, the MCF with $N_{\rm con}$ as the weight results in the amplitude and shape very comparable to those of the 2PCF, implying it is not statistically related to the local properties of each galaxy.
@@
Figure 7: Comparison of one-dimensional MCFs $s^{2}W_{0}(s)$ (top) and $W_{\Delta s}(\mu)$ (bottom) for four datasets: CMASS-NGC (red solid), COLA mocks with $\Omega_{m}=0.31$ (blue dashed-dotted), $\Omega_{m}=0.25$ (orange dashed), $\Omega_{m}=0.4$ (green dotted), respectively. From left to right, we present the results for the 2PCF and MCFs using the weights $N_{\rm con}$, $\bar{D}_{\rm nei}$, and $1/\bar{D}_{\rm nei}$, respectively. The grey-shaded region represents $2\sigma$ errors estimated using PATCHY mocks. For the COLA mock with $\Omega_{m}=0.31$, all derived MCFs in the $s$ and $\mu$ regions we considered ( $s\in[0,150]~{}h^{-1}\rm Mpc$, $\mu\in[0,0.97]$) match well with those of CMASS-NGC within $2\sigma$ level. However, when $\Omega_{m}=0.4$ or 0.25 in COLA largely deviates from the true value in CMASS, both 2PCFs and MCFs become significantly inconsistent with the observational data beyond $2\sigma$ level.
@@
Figure 8: Similar to Fig. 7, but for the normalized MCFs, $s^{2}\hat{W}_{0}(s)$ (top) in the range of $s\in[15,57]~{}h^{-1}{\rm Mpc}$, and $\hat{W}_{\Delta s}(\mu)$ (bottom) in the range of $\mu\in[0,0.97]$, equally spaced and divided into 7 bins. The grey-shaded region represents $2\sigma$ errors estimated by the PATCHY mocks. We observe that the results of COLA mocks under $\Omega_{m}=0.31$ agree well with those of CMASS-NGC, whereas larger and smaller $\Omega_{m}$ values can lead to apparent deviations, especially for the statistic $s^{2}\hat{W}_{0}$. Additionally, we note that the weight $1/\bar{D}_{\rm nei}$ can provide a sensitive probe to $\Omega_{m}$, as an incorrect value can result in significant deviations from those of the CMASS data beyond the $2\sigma$ level.
@@
Fig. 6 illustrates an analysis of the one-dimensional MCFs utilizing various weighting schemes on the CMASS-NGC dataset. On the left panel, contrasting with the 2PCF (red solid), the monopoles of MCFs, $s^{2}W_{0}(s)$ (as defined in Eq. 4), are illustrated for three distinct weights: $N_{\rm con}$ (blue dashed-dotted), $\bar{D}_{\rm nei}$ (orange dashed), and $1/\bar{D}_{\rm nei}$ (green dotted). The right panel depicts the same comparison but emphasizes the anisotropic clustering in MCFs, $W_{\Delta s}(\mu)$ (as defined in Eq. 5). Notably, the use of the weight $1/\bar{D}_{\rm nei}$ yields the highest magnitudes in both panels. This is attributed to the assignment of a larger weight to numerous objects in underdense regions, significantly enhancing the amplitudes of these MCFs. Conversely, assigning $\bar{D}_{\rm nei}$ as the weight results in the lowest amplitudes. Furthermore, in either panel, employing $N_{\rm con}$ as the weighting scheme for MCFs results in the amplitude and the shape that agree with those of the 2PCF. This is because $N_{\rm con}$ is not statistically related to the clustering strength of the structure (e.g., a galaxy in a low-density region can have a larger $N_{\rm con}$ than a galaxy in a high-density region). Therefore, using it as weights does not significantly alter the statistical clustering.
@@
In Fig. 7, a comparison of one-dimensional MCFs $s^{2}W_{0}(s)$ (top) and $W_{\Delta s}(\mu)$ (bottom) is presented for four datasets: CMASS-NGC (red solid), COLA mocks with $\Omega_{m}=0.31$ (blue dashed-dotted), $\Omega_{m}=0.25$ (orange dashed), and $\Omega_{m}=0.4$ (green dotted), respectively. The results for the 2PCF and MCFs using the weights $N_{\rm con}$, $\bar{D}_{\rm nei}$, and $1/\bar{D}_{\rm nei}$ are displayed from left to right. The grey-shaded region represents $2\sigma$ errors estimated using PATCHY mocks. For the COLA mock with $\Omega_{m}=0.31$, all derived MCFs in the $s$ and $\mu$ regions ($s\in[0,150],\mu\in[0,0.97]$) match well with those of CMASS-NGC within the $2\sigma$ level. However, when $\Omega_{m}=0.4$ or 0.25 in COLA, there is a significant deviation from the true value in CMASS, causing both 2PCFs and MCFs to become markedly inconsistent with the observational data beyond the $2\sigma$ level.
@@
Moreover, Fig. 8 illustrates the normalized MCFs, $s^{2}\hat{W}_{0}(s)$ (top) within the range $s\in[15,57]~{}h^{-1}{\rm Mpc}$, and $\hat{W}_{\Delta s}(\mu)$ (bottom) within the range $\mu\in[0.09,0.94]$, evenly spaced and divided into 7 bins. The grey-shaded region denotes $2\sigma$ errors estimated from PATCHY mocks. The results of COLA mocks under $\Omega_{m}=0.31$ are well consistent with those of CMASS-NGC. However, deviations become noticeable for larger and smaller $\Omega_{m}$ values, particularly for the statistic $s^{2}\hat{W}_{0}(s)$. Significantly, the weight $1/\bar{D}_{\rm nei}$ serves as a sensitive indicator for $\Omega_{m}$, where an incorrect value can lead to substantial deviations from those observed in the CMASS data, exceeding the $2\sigma$ level.
@@
The corresponding covariance matrix, denoted by $\bm{C}$, is evaluated using the PATCHY simulation mocks. Specifically, the empirical covariance matrix of a vector $\bm{p}$ is given by:
$\bm{C}$ $=$ $\left(\bm{p}-\bm{\bar{p}}\right)\left(\bm{p}-\bm{\bar{p}}\right)^{T}\right>$ (9)
@@
by averaging over all $N_{\rm mock}$ mock samples. Here, $\bm{p}_{i}$, with a length of $N_{p}$, denotes a vector containing all statistical quantities for the $i$-th mock sample. Note that the mean of $\bm{p}$ over all mocks is denoted by $\bm{\bar{p}}$. We utilized 1000 PATCHY mocks to accurately estimate the relevant covariance matrices for the various considered statistics.
@@
Figure 10: Changes of $\chi^{2}$ values (as defined in Eq. 7) for $\hat{W}_{0}(s)$ (left) and $\hat{W}_{\Delta s}(\mu)$ (right) as $\Omega_{m}$ varies from 0.25 to 0.4. Here, $\chi_{\rm min}^{2}$ is calculated by using $\Omega_{m}=0.31$ for $\chi^{2}$, where the $\chi^{2}$ values reach the minimum for either left and right panel compared to those for the other two $\Omega_{m}$ values. As seen, compared to the 2PCF, MCFs provide an apparently different dependence of $\chi^{2}$ on $\Omega_{m}$. As expected, employing a combination of 2PCF and MCFs with the three $\beta$-skeleton weights results in the highest sensitivity to $\Omega_{m}$. This is evident from the rapid changes in $\chi^{2}$ when $\Omega_{m}$ is varied.
@@
Displayed in Fig. 10 are the changes in $\chi^{2}$ values, according to Eq. 11, for $\hat{W}_{0}(s)$ (left) and $\hat{W}_{\Delta s}(\mu)$ (right) while $\Omega_{m}$ ranges from 0.25 to 0.4. The calculation of $\chi_{\rm min}^{2}$ utilizes the COLA mock with $\Omega_{m}=0.31$ following Eq. 7, where the minima in $\chi^{2}$ values are reached compared to those for the other two $\Omega_{m}$ values. As observed, MCFs exhibit a notably distinct dependence of $\chi^{2}$ on $\Omega_{m}$ compared to the 2PCF. As expected, the use of a combination of 2PCF and MCFs with the $\beta$-skeleton weights yields the highest sensitivity to $\Omega_{m}$. This is evident from the rapid changes in $\chi^{2}$ when varying $\Omega_{m}$.

### v3

@@ III.1 $\beta$-skeleton analysis
The deviations between the real data and COLA simulations, regardless of the $\Omega_{m}$ values used, are considerably small, with no significant differences observed. The mean comoving distance $r$ is 1346.1 $h^{-1}{\rm Mpc}$ for CMASS-NGC, 1345.7 $h^{-1}{\rm Mpc}$ for $\Omega_{m}=0.25$, 1346.7 $h^{-1}{\rm Mpc}$ for $\Omega_{m}=0.31$, and 1344.8 $h^{-1}{\rm Mpc}$ for $\Omega_{m}=0.4$. These values suggest that when varying $\Omega_{m}$, the changes in $r$ for COLA are on the order of 0.1%, making it challenging to rely solely on this quantity for distinguishing between cosmologies.
Similarly, the density distributions of the connection length $L_{\rm con}$ and the orientation of these connections, $\mu$, are illustrated in the middle and right panels. We observe that the discrepancies in either $L_{\rm con}$ or $\mu$ between real data and COLA simulations are statistically insignificant. For instance, the mean $L_{\rm con}$ is 8.83 ${}h^{-1}{\rm Mpc}$ for CMASS-NGC, and 8.70, 8.68, and 8.71 $h^{-1}{\rm Mpc}$ for $\Omega_{m}=0.25,0.31,$ and $0.4$, respectively, leading to deviations of less than $2\%$ across all cases. Additionally, the mean values of $\mu$, averaged over the range of $\mu\in[0,0.97]$, are 0.50 for CMASS-NGC, and 0.51, 0.52, 0.53 for $\Omega_{m}=0.25,0.31,0.4$ in COLA, respectively. This suggests that the directions of the connections in all cases are consistently randomly distributed, with no preferred orientation. Moreover, we can clearly observe a rise in density when $\mu>0.9$. This is likely due to the Finger of God (FoG) effect, increasing the orientation along LOS.
Figure 4: Histogram of the number of neighbors around each galaxy, referred to as connectivity ($N_{\rm con}$), in the range of $[0,5]$ within the $\beta$-skeleton cosmic webs. It peaks at $N_{\rm con}=2$ and then decreases rapidly. The variation in connectivity is negligible for both CMASS-NGC data and COLA mocks under three different $\Omega_{m}$ values.
@@
On the right side, four violin plots depict the dependence of $\bar{D}_{\rm nei}$ distributions on connectivity $N_{\rm con}$, ranging from 1 to 5. The cases of CMASS-NGC data and COLA mocks under three different $\Omega_{m}$ values are presented sequentially. As seen, all COLA results are nearly identical to the CMASS-NGC case when $N_{\rm con}<5$. However, when $N_{\rm con}=5$, the distribution of $\bar{D}_{\rm nei}$ exhibits slight variation, with mean values of $\bar{D}_{\rm nei}=11.52,12.38,11.27,12.78~{}h^{-1}{\rm Mpc}$ for CMASS-NGC and COLA simulations with $\Omega_{m}=0.31$, $0.25$, and $0.4$.
@@ III.2 MCFs using the $\beta$-skeleton statistics
Briefly, the procedure for calculating the MCFs remains the same as the standard one; however, the weights are assigned based on the measures of the $\beta$-skeleton statistics. Compared with the traditional 2PCF, defined as $\xi(\bm{r})=\langle\delta(\bm{x})\delta(\bm{x}+\bm{r})\rangle$, the form of MCF, following Yang et al. (2020), is given by
@@
where the term $w(\bm{x})$ denotes the chosen weights used in MCFs. The term $\delta(\bm{x})$ denotes the pointlike density contrast, expressed as $\delta(\bm{x})=\delta{\rho}/\bar{\rho}$.
@@
We utilize the widely adopted Landy–Szalay estimator Landy and Szalay (1993), expressed as
@@
The random sample comprises 10 times the number of objects compared to the CMASS-NGC galaxies, and each PATCHY simulation for covariance estimation includes 20 times the number of objects than CMASS-NGC. Based on our testing, results converge with a random sample containing about 10 times more objects than CMASS-NGC, and the weights for the random samples are consistently set to unity.
@@ III.3 One-dimensional MCFs: $\hat{W}_{0}(s)$ and $\hat{W}_{\Delta s}(\mu)$
By integrating $W(s,\mu)$ over either $s$ or $\mu$, two one-dimensional statistical quantities can be defined. The first one is the monopole of the MCF, a function of the clustering scale, represented as
@@
The second quantity is the $\mu$-dependent function, represented as
@@
To mitigate the impact of galaxy bias and enhance the accuracy of the analysis, it is common to employ normalized statistics, which utilize the shape rather than the amplitude to extract cosmological information. By doing this, the normalized quantities based on Eqs. (4) and (5) can be expressed as follows:
@@
In choosing $a=15~{}h^{-1}{\rm Mpc}$, $b=57~{}h^{-1}{\rm Mpc}$, and $\mu_{\max}=0.97$ (to reduce effects of fiber collisions and RSDs), extensive testing indicates that the selected parameters are effective for studying clustering features and for enhancing the sensitivity on different cosmologies. From tests, we find that opting for equally spaced divisions of $s$ within the range of $s\in[15,57]~{}h^{-1}{\rm Mpc}$ and $\mu$ in the interval $\mu\in[0,0.97]$ into seven bins results in optimal and robust constraints on $\Omega_{m}$.
@@ IV Results
Figure 6: Comparison of one-dimensional MCFs with different weighting schemes for the CMASS-NGC dataset. Left: In comparison to the 2PCF (red solid), the monopole of MCFs, $s^{2}W_{0}(s)$ [as defined in Eq. (4)], is shown for three different weights: $N_{\rm con}$ (blue dashed-dotted), $\bar{D}_{\rm nei}$(orange dashed), and $1/\bar{D}_{\rm nei}$ (green dotted). Right: Same as in the left panel, but for focusing on the anisotropic clustering in MCFs, $W_{\Delta s}(\mu)$ [as defined in Eq. (5)]. As observed, using the weight $1/\bar{D}_{\rm nei}$ leads to the highest magnitudes in both panels. This is because a number of objects in underdense regions are assigned a larger weight, significantly enhancing the amplitudes of these two-point functions. Conversely, assigning $\bar{D}_{\rm nei}$ as the weight will lead to the lowest amplitudes in both panels. In either the left or right panel, the MCF with $N_{\rm con}$ as the weight results in the amplitude and shape very comparable to those of the 2PCF, implying it is not statistically related to the local properties of each galaxy.
@@
Figure 7: Comparison of one-dimensional MCFs $s^{2}W_{0}(s)$ (top) and $W_{\Delta s}(\mu)$ (bottom) for four datasets: CMASS-NGC (red solid), COLA mocks with $\Omega_{m}=0.31$ (blue dashed-dotted), $\Omega_{m}=0.25$ (orange dashed), $\Omega_{m}=0.4$ (green dotted), respectively. From left to right, we present the results for the 2PCF and MCFs using the weights $N_{\rm con}$, $\bar{D}_{\rm nei}$, and $1/\bar{D}_{\rm nei}$, respectively. The gray-shaded region represents $2\sigma$ errors estimated using PATCHY mocks. For the COLA mock with $\Omega_{m}=0.31$, all derived MCFs in the $s$ and $\mu$ regions we considered ($s\in[0,150]~{}h^{-1}\rm Mpc$, $\mu\in[0,0.97]$) match well with those of CMASS-NGC within $2\sigma$ level. However, when $\Omega_{m}=0.4$ or 0.25 in COLA largely deviates from the true value in CMASS-NGC, both 2PCFs and MCFs become significantly inconsistent with the observational data beyond $2\sigma$ level.
@@
Figure 8: Similar to Fig. 7, but for the normalized MCFs, $s^{2}\hat{W}_{0}(s)$ (top) in the range of $s\in[15,57]~{}h^{-1}{\rm Mpc}$, and $\hat{W}_{\Delta s}(\mu)$ (bottom) in the range of $\mu\in[0,0.97]$, equally spaced and divided into 7 bins. The grey-shaded region represents $2\sigma$ errors estimated by the PATCHY mocks. We observe that the results of COLA mocks under $\Omega_{m}=0.31$ agree well with those of CMASS-NGC, whereas larger and smaller $\Omega_{m}$ values can lead to apparent deviations, especially for the statistic $s^{2}\hat{W}_{0}$. Additionally, we note that the weight $1/\bar{D}_{\rm nei}$ can provide a sensitive probe to $\Omega_{m}$, as an incorrect value can result in significant deviations from those of the CMASS-NGC data beyond the $2\sigma$ level.
@@
Figure 6 illustrates an analysis of the one-dimensional MCFs utilizing various weighting schemes on the CMASS-NGC dataset. On the left panel, contrasting with the 2PCF (red solid), the monopoles of MCFs, $s^{2}W_{0}(s)$ [as defined in Eq. (4)], are illustrated for three distinct weights: $N_{\rm con}$ (blue dashed-dotted), $\bar{D}_{\rm nei}$ (orange dashed), and $1/\bar{D}_{\rm nei}$ (green dotted). The right panel depicts the same comparison but emphasizes the anisotropic clustering in MCFs, $W_{\Delta s}(\mu)$ [as defined in Eq. (5)]. Notably, the use of the weight $1/\bar{D}_{\rm nei}$ yields the highest magnitudes in both panels. This is attributed to the assignment of a larger weight to numerous objects in underdense regions, significantly enhancing the amplitudes of these MCFs. Conversely, assigning $\bar{D}_{\rm nei}$ as the weight results in the lowest amplitudes. Furthermore, in either panel, employing $N_{\rm con}$ as the weighting scheme for MCFs results in the amplitude and the shape that agree with those of the 2PCF. This is because $N_{\rm con}$ is not statistically related to the clustering strength of the structure (e.g., a galaxy in a low-density region can have a larger $N_{\rm con}$ than a galaxy in a high-density region). Therefore, using it as weights does not significantly alter the statistical clustering.
@@
In Fig. 7, a comparison of one-dimensional MCFs $s^{2}W_{0}(s)$ (top) and $W_{\Delta s}(\mu)$ (bottom) is presented for four datasets: CMASS-NGC (red solid), COLA mocks with $\Omega_{m}=0.31$ (blue dashed-dotted), $\Omega_{m}=0.25$ (orange dashed), and $\Omega_{m}=0.4$ (green dotted), respectively. The results for the 2PCF and MCFs using the weights $N_{\rm con}$, $\bar{D}_{\rm nei}$, and $1/\bar{D}_{\rm nei}$ are displayed from left to right. The gray-shaded region represents $2\sigma$ errors estimated using PATCHY mocks. For the COLA mock with $\Omega_{m}=0.31$, all derived MCFs in the $s$ and $\mu$ regions ($s\in[0,150],\mu\in[0,0.97]$) match well with those of CMASS-NGC within the $2\sigma$ level. However, when $\Omega_{m}=0.4$ or 0.25 in COLA, there is a significant deviation from the true value in CMASS-NGC, causing both 2PCFs and MCFs to become markedly inconsistent with the observational data beyond the $2\sigma$ level.
@@
Moreover, Fig. 8 illustrates the normalized MCFs, $s^{2}\hat{W}_{0}(s)$ (top) within the range $s\in[15,57]~{}h^{-1}{\rm Mpc}$, and $\hat{W}_{\Delta s}(\mu)$ (bottom) within the range $\mu\in[0,0.97]$, evenly spaced and divided into seven bins. The gray-shaded region denotes $2\sigma$ errors estimated from PATCHY mocks. The results of COLA mocks under $\Omega_{m}=0.31$ are well consistent with those of CMASS-NGC. However, deviations become noticeable for larger and smaller $\Omega_{m}$ values, particularly for the statistic $s^{2}\hat{W}_{0}(s)$. Significantly, the weight $1/\bar{D}_{\rm nei}$ serves as a sensitive indicator for $\Omega_{m}$, where an incorrect value can lead to substantial deviations from those observed in the CMASS-NGC data, exceeding the $2\sigma$ level.
@@
The corresponding covariance matrix, denoted by $\bm{C}$, is evaluated using the PATCHY simulation mocks. Specifically, the empirical covariance matrix of a vector $\bm{p}$ is given by
$\bm{C}$ $=$ ⟨$\left(\bm{p}-\bm{\bar{p}}\right)\left(\bm{p}-\bm{\bar{p}}\right)^{T}\right>$ (9)
@@
by averaging over all $N_{\rm mock}$ mock samples. Here, $\bm{p}_{i}$, with a length of $N_{p}$, denotes a vector containing all statistical quantities for the $i$th mock sample. Note that the mean of $\bm{p}$ over all mocks is denoted by $\bm{\bar{p}}$. We utilized 1000 PATCHY mocks to accurately estimate the relevant covariance matrices for the various considered statistics.
@@
Figure 10: Changes of $\chi^{2}$ values [as defined in Eq. (7)] for $\hat{W}_{0}(s)$ (left) and $\hat{W}_{\Delta s}(\mu)$ (right) as $\Omega_{m}$ varies from 0.25 to 0.4. Here, $\chi_{\rm min}^{2}$ is calculated by using $\Omega_{m}=0.31$ for $\chi^{2}$, where the $\chi^{2}$ values reach the minimum for either left and right panel compared to those for the other two $\Omega_{m}$ values. As seen, compared to the 2PCF, MCFs provide an apparently different dependence of $\chi^{2}$ on $\Omega_{m}$. As expected, employing a combination of 2PCF and MCFs with the three $\beta$-skeleton weights results in the highest sensitivity to $\Omega_{m}$. This is evident from the rapid changes in $\chi^{2}$ when $\Omega_{m}$ is varied.
@@
Displayed in Fig. 10 are the changes in $\chi^{2}$ values, according to Eq. (11), for $\hat{W}_{0}(s)$ (left) and $\hat{W}_{\Delta s}(\mu)$ (right) while $\Omega_{m}$ ranges from 0.25 to 0.4. The calculation of $\chi_{\rm min}^{2}$ utilizes the COLA mock with $\Omega_{m}=0.31$ following Eq. (7), where the minima in $\chi^{2}$ values are reached compared to those for the other two $\Omega_{m}$ values. As observed, MCFs exhibit a notably distinct dependence of $\chi^{2}$ on $\Omega_{m}$ compared to the 2PCF. As expected, the use of a combination of 2PCF and MCFs with the $\beta$-skeleton weights yields the highest sensitivity to $\Omega_{m}$. This is evident from the rapid changes in $\chi^{2}$ when varying $\Omega_{m}$.